OpenStudy (anonymous):

Anyone help me with this question and the first person who does ill give a medal

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

The area of a rectangle can be found by multiplying the length by the width. A rectangular postage stamp has a length of nine-tenths inch and a width of five-sixths inch. What is the area of the stamp, in square inches? three-fourths square inches one-seventeenth square inches one-half square inches one and one-fourth square inches

OpenStudy (kc_kennylau):

Area:9/10*5/6=45/60=3/4 (A)

OpenStudy (anonymous):

kc is right

OpenStudy (kc_kennylau):

thx